# A new economic and financial theory of money

### v5

@@ XII Example of an electronic currency at work
XIII Government and taxation
@@
We now turn our attention to the role of government and taxation. The first role of government is to make the rules of the economic and social game, that is law, and to enforce them. It is important that the rules be kept to a minimum and that the rules respect the freedoms discussed in Sec. XIV. Government must also keep the records on property rights. In addition to these tasks, Government is responsible for the common physical and social infrastructure. This includes: (1) defense, (2) social security, (3) education, (4) arts, (5) public infrastructure (such as roads, railroads, and parks), (6) healthcare, and (7) fundamental and government related research.
@@
Examples of rules that are conservative are transaction & sales taxes and equity & electronic currency funding which have repayment/issue terms that are conservative forces
@@
Examples of rules that are not conservative are wealth & property taxes and loans with terms of debt $\Delta R=-\nu R\Delta\tau$ or $R(q,\tau)=R(q)\text{e}^{-\nu\tau}$.
@@
XIV Freedom of labor, intellectual property, and law
We start by discussing the concept of “Communism & Freedom”. Milton Friedman went out of his way in the early 1960’s to cast capitalism as being superior to communism, but he made the point that it was essential that there be freedom so that the private sector could zealously pursue its maximization of profit. If it was profitable, then it was in the social good. This is fundamentally flawed. Greed is not good, it is evil. Friedman’s ideas formed the basis of Reagan and Thatcher conservatism of the 1980’s.
@@ XIV Freedom of labor, intellectual property, and law
It should be noted that the combination of “Communism & Freedom” that is being proposed is more economically conservative, more truly free or privatized, than existing «Communisme et Liberté»  in free societies such as France. There is no one central government control that has a “Robinhood” policy of taking from the rich and giving to the poor, with an inescapable central government planning. Instead there are multiple liberated private sub-economies that significantly raise the virtuous economic activity with some of the increase going to the wealthy, modestly increasing their wealth, but a much larger part of the increase going to the poor, significantly increasing their wealth and bringing them solidly into the middle class.
@@
Protection comes from the insular nature of the sub-economy, and the ability of the leader of the sub-economy to directly support (subsidize) the economic activity, effectively acting as a protective import duty. For example, a leader of a sub-economy provides salaries in the currency of the sub-economy for artists based on how much of their art is consumed by the sub-economy. The same is true of creators of intellectual property. As discussed in Sec. XIII, some of this compensation and employment may be by Government Inc. with the revenue coming from a transactional tax. Tokens of ownership (e.g., NFTs) or certification of action (e.g., carbon credits) are just digital certificates and should have nothing to do with a currency.
@@
Many of these freedoms also eliminate dissipation from the economic system which maintains the conservative dynamics of the economy, as discussed in Sec. XIII.
XV Resilience of economic system to economic collapse
@@
Things are not much better with crypto currencies. Proof-of-work crypto currencies, like Bitcoin, rely on a mining algorithm to control the creation of the currency. There is no mechanism to reduce the money supply. The creation of the currency is at a rate that varies depending on the time it takes to solve the elliptic problem on a computer. Given that the economic system response has multiple time scales, and there will be times when the money supply should be decreased, this is again a poor system of system control. The situation is even worse for proof-of-stake systems like the New Ethereum. They have built-in, through the integration and debt-like coupon payments, a phase shift that leads to system instability. There is also the large financial leverage that is built into the financial structure of these crypto currencies because of lack of both liquid and capital reserves, that will be discussed in Sec. XVI and is shown graphically in Fig. 22. This can and has led to large volatility in the value of crypto currencies and their economic collapse. There is rarely any mechanism for fiscal investment programmed into the crypto currencies. When it is programmed into the currencies, it is not based on a rigorous economic evaluation of the investment opportunities.
@@
The electronic currency that we are proposing is very different. First of all it has significant levels of liquid assets and both tangible and intangible capital assets, as shown in Fig. 4. This provides a financial heat bath to give inertia to the electronic currency that will lead to a very stable value for the electronic currency. In addition, the electronic currency management firm will use an advanced AI control system, based on the multi-scale system response of the sub-economy, that will further reduce the fluctuations in the value of the currency and stabilize the economy. The currency management firm will coordinate the fiscal investment policy with the monetary policy, issuing the currency needed for investment based on the projected demand that those investments will create. This will lead to a modest long-term stable growth in the value of the currency.
The reserves are essential to the stability of the economy. They give the dynamical financial system the inertia to change that stabilizes the economy. As shown in Fig. 4, the financial structure of the currency equity (fully backed security by tangible assets) and the stock equity (fully backed security by intangible assets) eliminates the super-charged risk of a highly leveraged derivative security as shown in Fig. 22. First of all, the liquid reserves (whether that be gold, cash or inventory like stored hydrocarbons) give the assets that can be sold by the currency management firm to purchase the electronic currency as directed by the AI, to stabilize the value of the electronic currency. It is anticipated that less cash reserves are needed for a larger sub-economy since the relative size of the fluctuations are smaller. Then there is the longer-term strategic monetary policy associated with execution of the fiscal investment policy and the control of longer term inflation and deflation. In times of deflation (recession), this will involve issuance of new electronic currency to buy liquid assets, subsidize labor costs, buy capital assets, and invest in existing capital assets. This will position the currency management firm to respond to periods of inflation by selling the liquid assets, terminating the labor subsidies, selling capital assets and reducing the investment in existing capital assets. This buying and selling is in contrast to the loaning and borrowing of sovereign currencies. Most governments do not have either the liquid assets (even if they do have a gold reserve like the US had at one time they do not actively trade it, it is only there to be a physical standard), or the capital assets to sell in a time of inflation. They rely only on the issuance of debt. A notable exception is the recent sale of the US petroleum reserve to combat the post COVID inflation. There are additional ways that the reserves increase the resiliency that will be discussed in the following paragraphs of this section.
@@
Before we proceed to that phase, let us examine why the system that we are proposing does not have a “risk squeeze”. The reason is simple. This system does not make the local assumption and securitizes the loan using the real model of risk – a multi-scale model of risk using AI.
@@
In summary, our system is protected from economic collapse by: (1) significant (about 10%) liquid reserves, (2) complete capital reserves (that is the electronic currency being used for transactions and working capital), (3) AI and non-debt based methods of monetary control, with insurance, matched to the structure of the economy, (4) a floating currency value, and (5) the use of electronic currency leveraged, not debt leveraged financing.
XVI Existing currencies
Existing currencies are a grab bag of misfits that range from the mundane to fraud. Let’s start by examining the ubiquitous sovereign currency, that has been the only type of currency until recently. Historical currencies have been based on a precious metal (gold or silver in most cases) liquid reserve that has literally been attached as a standard to the coin as the material from which it was fabricated. In more recent history, paper notes were issued and pledged that there were gold or silver reserves in the vault of the sovereign. In the most recent history, sovereigns have floated the currency, eliminating the gold reserves and backing the currency with the tangible and intangible assets of the sovereignty, with negligible liquid reserves in many cases. Monetary policy has been controlled by the literal minting of coins and printing of currency, then having the central bank loan to commercial banks at a prime interest rate. To reduce the currency in circulation the sovereign issues debt (i.e., US T-bills and British GILTs). Rarely has the sovereign bought and sold liquid reserves as an method of implementing monetary policy. A recent exception is the selling of the US Petroleum Reserve by the US Government as a measure to combat the post-COVID supply-chain induced inflation. As discussed previously in Sec. XV, this is a poor method of system control at best and is unstable in many cases. In addition, there is usually an uncoordinated fiscal investment policy that is driven by politics, not economics.

### v6

@@ XII Example of an electronic currency at work
XIII Ubuntu business model
The example presented in Sec. XII deploys a new business model based on transactional equity. This Ubuntu business model removes the dissipation that is embedded in the conventional debt based business model that uses borrow/loan financing – creating an economic superconductor free of resistance, that is debt. Although the dissipation, that is interest, controls and stabilizes the economic system, it places a large constraint on the operations of the business reducing the revenue by 90% or more, and robs the business of free cash flow needed for its sustenance and growth. The dissipation does dominate the business performance forecasting, resulting in a simple exponential random walk or drift-diffusion to zero that is the solution to the easily solved Black-Scholes or Fokker-Planck equation. In contrast, when the resistance is removed by use of the Ubuntu business model and replaced by genAI based economic control (forecasting for investment) and stabilization, the economy is put into motion by investment then is self sustaining without need for further investment, like a superconductor.
The Ubuntu business model recognizes that the economy is a collective of individuals, as such, there emerges a collective behavior or virtual individual that determines, that is controls, the collective behavior. This control emanates from a simply beautiful symmetry – the coordination of the economically interacting individuals creating an economic force, and the economic force then coordinating the economic trade of the individuals. Another way of looking at this is that the actions of the virtual individual, who follows a geodesic motion, are approximated by genAI. The genAI then generates and multiply reflects the actions of the virtual individual through a Hall of Mirrors to obtain, that is forecast, the economic evolution of the collective. This is demonstrated by “The Mirror Maze” scene from the 1928 Charlie Chaplin movie “The Circus” (Chaplin, 1928) and shown in Fig. 22.
Figure 22: Demonstration of how the motion of the virtual individual is multiply reflected by a Hall of Mirrors to give the economic evolution of the collective. This is shown in the “The Mirror Maze” scene from the 1928 Charlie Chaplin movie “The Circus” ( YouTube video).
Given how the actions of the virtual individual are multiply reflected through the Hall of Mirrors, the state of economic prosperity can be stabilized by genAI ponderomotive stabilization. This method is similar to the way that a sheepdog herds sheep. The sheepdog runs around the herd very fast, compared to movements of the herd, nipping at the heals of the sheep (vibrating them) if they wonder away from the mountain pass, that is the state of economic prosperity. The sheepdog is effectively creating a small alpine valley at the mountain pass. Practically, this is a high frequency arbitrage trading of the transactional equity that provides liquidity.
We use the name Ubuntu (ubu, 2024; Mandela, 2004) to describe the business model, the new unified economic theory, and the genAI because all three are based on the conservative interactions of collective members, that is the “interconnectedness” of the Spirit of Ubuntu. The Spirit of Ubuntu can be viewed as the virtual individual of the collective. We believe that we are “bringing back the African Spirit of Ubuntu” with this business model.
The Ubuntu way of “mutuality” is based on the concept of sharing, putting toys in a community toy chest so that there are more toys from which to choose as shown in Fig. 23. The community software chest is GitHub. The community capital chest is Martin Luther’s Community Chest of the Protestant Reformation (Luther, 1523). The electronic transaction equity version of community capital chest is Ubuntu Financing. This is related to the Economics of Mutuality (Mayer and Roche, 2021), Modern Monetary Theory (Kelton, 2020), and Special Purpose Acquisition Companies (SPACs).
Figure 23: The good of sharing. (a) Children putting toys in a community toy chest so that there are more toys from which to choose, leading to happiness. (b) Children being greedy, keeping their toys to themselves, leading to sadness.
When the Ubuntu business model eliminates the dissipation from the business by the use of Ubuntu (print/invest) Financing with transactional equity, another method of control and forecasting must be found. This control and forecasting is what is provided by the new Ubuntu genAI. The difference between drift-diffusive forecasting and genAI forecasting is shown in Fig. 3. The resistive control, that is going “full gas” and applying the brakes to change the speed shown in Fig. 24, is replaced by the ponderomotive genAI control discussed in Sec. VII and shown in Fig. 7. The result of removing the dissipation without implementing ponderomotive control, the effect of resistive control, and the economic Nirvana created by ponderomotive genAI control are shown in Fig. 25.
Figure 24: Illustration of resistive control. Going “full gas” on a motorcycle and applying the brakes to change the speed. Image generated by OpenAI’s DALL-E.
Figure 25: Illustrations of different types of control. (a) Jimmy Carter tries to shepherd the new free society (that is, The Great Society) emerging from the 1960’s, which wanders out of control. (b) The dystopian response is Lewis Powell and the cabal locking up society. (c) Because of transactional equity and genAI, we are at the dawn of a new era of sustainable economic prosperity, in an alpine valley shepherded by genAI. Images generated by OpenAI’s DALL-E.
The Ubuntu business model is characterized by: (1) maximization of sustainable revenue, (2) business cycle (2 to 6 years) of transactional equity savings and inventory (both input and output), (3) forecast and control using genAI models, (4) Ubuntu (print/invest) Financing, (5) stakeholders (employees, suppliers and customers) being stockholders (owners), and (5) social responsibility (respecting the dignity of employees, suppliers and customers). In contrast, the conventional debt based business model is characterized by: (1) maximization of short term profit, (2) transaction cycle (15 days or Just-In-Time) of cash savings and inventory, (2) forecast and control using using diffusive models (i.e., Discounted Cash Flow analysis or the Black-Scholes Equation for options and derivatives pricing), (3) debt based (borrow/loan) financing, (4) stockholders not being stakeholders, and (5) exploitation of society (that is, employees, suppliers, customers, and intellectual property).
The conventional debt based business model, without sufficient savings, has poor resiliency and performance. Production, that is revenue, is constrained so that there is not full employment, supply is limited, and prices are artificially high to ration supply. The economic collective is constrained so that there is never surplus supply, that is all production is immediately sold. In contrast, the Ubuntu business model generates 10 times the revenue or value, makes 100 times the investment, has plentiful supply, is resilient to economic shocks, and has low prices.
Ubuntu Financing is based on the common stock equity and transactional equity model shown in Fig. 4. Transactional equity primarily securitizes the low risk tangible assets of the business, while the common stock equity securitizes the high risk intangible assets (things like value-in-place, goodwill, brand value, know-how, show-how, and intellectual property). The risk (fluctuations) of the transactional equity are further reduced by the ponderomotive stabilization which transfers risk from the transactional equity to the common stock equity. Such transactional equities are ubiquitous, whether that be Zelle, Venmo, casino chips, airline miles, StarBucks, gift cards, loyalty programs, subway tokens, REI dividends, or carnival tickets.
The Ubuntu firm will operate the Private Equity (PE) fund of the future. The holders of the common stock equity take the place of the limited partners of the PE fund. Conventional debt leveraging is replaced by Ubuntu Financing, that is the printing of transactional equity and the subsequent acquisitions or investments with that transactional equity. This is equivalent to an all stock acquisition, for instance when Exxon printed \$60 billion in stock then bought Pioneer Resources with that stock. The only difference with the Ubuntu firm is that one can tap-and-pay with transactional equity. One does not first need to exchange the common stock for coin. Transactional equity is a centrally accounted (not on a distributed ledger like the blockchain, that is it is not crypto) electronic coin. This PE fund of the future does not have meddlesome limited partners, does not have banks that require regular coupon payments, can make much larger investments, and generates much more return, that is value.
XIV Government and taxation
@@
We now turn our attention to the role of government and taxation. The first role of government is to make the rules of the economic and social game, that is law, and to enforce them. It is important that the rules be kept to a minimum and that the rules respect the freedoms discussed in Sec. XV. Government must also keep the records on property rights. In addition to these tasks, Government is responsible for the common physical and social infrastructure. This includes: (1) defense, (2) social security, (3) education, (4) arts, (5) public infrastructure (such as roads, railroads, and parks), (6) healthcare, and (7) fundamental and government related research.
@@
Examples of rules that are conservative are transaction and sales taxes, and equity and electronic currency funding which have repayment/issue terms that are conservative forces
@@
Examples of rules that are not conservative are wealth and property taxes, and loans with terms of debt $\Delta R=-\nu R\Delta\tau$ or $R(q,\tau)=R(q)\text{e}^{-\nu\tau}$.
@@
XV Freedom of labor, intellectual property, and law
We start by discussing the concept of “Communism & Freedom”. Milton Friedman went out of his way in the early 1960’s to cast capitalism as being superior to communism, but he made the point that it was essential that there be freedom so that the private sector could zealously pursue its maximization of profit. If it was profitable, then it was in the social good. This is fundamentally flawed. Greed is not good, it is evil. What is profitable is exploiting society. The resistivity that the pursuit of profitability imposes on society, leads to oppression that opposes freedom (Glinsky, 2024b). Friedman’s ideas formed the basis of Reagan and Thatcher conservatism of the 1980’s.
@@ XV Freedom of labor, intellectual property, and law
It should be noted that the combination of “Communism & Freedom” that is being proposed is more economically conservative, more truly free or privatized, than existing «Communisme et Liberté»  in free societies such as France. There is no one central government control that has a “Robinhood” policy of taking from the rich and giving to the poor, with an inescapable central government planning. Instead there are multiple liberated private sub-economies that significantly raise the virtuous economic activity with some of the increase going to the wealthy, modestly increasing their wealth, but a much larger part of the increase going to the poor, significantly increasing their wealth and bringing them solidly into the middle class. As we will discuss at the end of this section, this is a Democratic Socialism, not the centrally controlled Autocratic Socialism that Friedman criticized.
@@
Protection comes from the insular nature of the sub-economy, and the ability of the leader of the sub-economy to directly support (subsidize) the economic activity, effectively acting as a protective import duty. For example, a leader of a sub-economy provides salaries in the currency of the sub-economy for artists based on how much of their art is consumed by the sub-economy. The same is true of creators of intellectual property. As discussed in Sec. XIV, some of this compensation and employment may be by Government Inc. with the revenue coming from a transactional tax. Tokens of ownership (e.g., NFTs) or certification of action (e.g., carbon credits) are just digital certificates and should have nothing to do with a currency.
@@
Many of these freedoms also eliminate dissipation from the economic system which maintains the conservative dynamics of the economy, as discussed in Sec. XIV.
The issue of freedom can be understood vis-a-vis the interaction of two things. The first thing is the objective function that is optimized, whether that is DCF or GDP. This is equivalent to the method of control, whether that is genAI or resistivity. It is whether socialism is controlling the economic collective, or it is capitalism. Whether it is controlled via profit maximizing capitalism, or it is controlled with revenue maximizing socialism.
The second thing is the distribution of control. Whether it is a distributed web of economic communes or a single economic commune centrally controlled by a Central Bank and a centralized government or Central Committee. This is either autocracy or democracy. It is whether you end up with monopolies, both economic and political, or free-markets.
Countries can now be put into quadrants of this matrix, shown in Fig. 26. First, there are the Socialist Autocracies like USSR, China, Cuba, and North Korea. Second, there are the Socialist Democracies like the Scandinavian countries, to a lesser extent, France, Portugal, Australia, Canada, Switzerland, and Germany. It is very interesting that most indigenous cultures of Africa, Americas and Australia are Socialist Democracies. Third, there are the Capitalist Democracies like United States of the 1960’s, initially Israel, Hungary, Turkey, and Russia underneath Gorbachev and Yeltsin. Finally the Capitalist Autocracies like Hitler’s Third Reich, Putin’s Russia, Trump’s USA, Erdoğan’s Turkey, Orbán’s Hungary, Netanyahu’s vision for Israel, and essentially most of the history of Western Civilization. Note that Capitalist Autocracies are stable with poor economic performance. Socialist Autocracies are unstable, mostly because of their very poor performance. Capitalistic Democracies are also unstable with reasonable or modest performance. The best and stable economies are Socialistic Democracies.
Figure 26: Matrix of economic collective systems based on the objective and distribution of control.
We now go through these types of economic collective systems again. Democratic Socialism is stable with high performance. They are rated the happiest countries in the world. This was successfully deployed in the American West from 1830 to 1860 (Galbraith, 2017). It was developed using multiple paper currencies, led by the Wells Fargo Bank, founded in 1852. Each bank covered one day’s horse ride and had its own promissory note or currency that it could print and grant. During this time there was a two times decrease in prices due to increased productivity, that is deflation. There was a 100x increase in GDP of the American West over this time. This movement continued after the US Civil War, as the Greenback Party (1874-1889). A good example of an economic cooperative, that exists today, is the REI Co-op for recreational equipment. It is 90 years old and counting. It is the premier US retailer of camping gear. It currently has \$4 billion per year in revenue and 15,000 extremely happy employees. This leads to a distribution of wealth, power, and control that customizes the guidance, that is sharing, leading to and reinforcing a free market and democracy. In contrast, there is Autocratic Capitalism that is stable with poor performance or Autocratic Socialism which is unstable with very poor performance. The much studied one is Democratic Capitalism, which is familiar to most economics departments and business schools. It is the capitalistic Monetarism (“Capitalism & Freedom”) of Milton Friedman and the University of Chicago school of economics. Note that commercial firms, under this method of control, have very short lifetimes, an average of about 15 years and shrinking. This is because the diffusive model does not capture disruptions (does not model them), and the force of profit prevents the transition of firms through market disruptions. Old firms, like Sears, fail and are replaced by new firms, like Amazon, when there is a market disruption. The resistivity of 6.7%, embedded in the economy via a typical AAA bond rate, constrains and grinds businesses to a halt in about 15 years. The lifetime of the business is one over the interest rate, $t=1/\nu$. The imposition of the interest rate is a forecast that businesses will go to zero revenue in 15 years, but imposition of the interest causes the business to have zero revenue in 15 years. It is a self-fulfilling prophesy. There is a strong natural force to concentrate wealth, power, control and guidance, so that there will be an evolution to autocracy and monopoly. Greed will be an irresistible force, despite regulations and democratic constitutional constraints, that will prevail resulting in monopolies and autocracies despite one’s best efforts to prevent them.
The current lifespan of a business in the US (that is, a S&P 500 company) is 18 years and shrinking, down from 35 years in 1965. Only 50 companies are still on the Fortune 500 that were there 70 years ago. With respect to the Dow Jones index, only one company (GE) is still on the index that was there in 1896 when it was formed. There are four companies added in the late 1920’s and 1930’s (ExxonMobil 1928, Procter & Gamble 1932, DuPont 1935, and United Technologies 1939). The next oldest is 3M added in 1976. Economists call this “creative destruction” that fuels economic prosperity. Isn’t it rather corporate destruction driven by profit that destroys economic prosperity? No business, other than IBM who has done it twice, has been able to survive a major disruption to their business model. It is interesting that a lifetime of 15 years corresponds to a discount rate of 6.7%. This is not a coincidence.
XVI Resilience of economic system to economic collapse
@@
Things are not much better with crypto currencies. Proof-of-work crypto currencies, like Bitcoin, rely on a mining algorithm to control the creation of the currency. There is no mechanism to reduce the money supply. The creation of the currency is at a rate that varies depending on the time it takes to solve the elliptic problem on a computer. Given that the economic system response has multiple time scales, and there will be times when the money supply should be decreased, this is again a poor system of system control. The situation is even worse for proof-of-stake systems like the New Ethereum. They have built-in, through the integration and debt-like coupon payments, a phase shift that leads to system instability. There is also the large financial leverage that is built into the financial structure of these crypto currencies because of lack of both liquid and capital reserves, that will be discussed in Sec. XVII and is shown graphically in Fig. 27. This can and has led to large volatility in the value of crypto currencies and their economic collapse. There is rarely any mechanism for fiscal investment programmed into the crypto currencies. When it is programmed into the currencies, it is not based on a rigorous economic evaluation of the investment opportunities.
@@
The electronic currency that we are proposing is very different. First of all it has significant levels of liquid assets and both tangible and intangible capital assets, as shown in Fig. 4. This provides a financial heat bath to give inertia to the electronic currency that will lead to a very stable value for the electronic currency. In addition, the electronic currency management firm will use an advanced genAI control system, based on the multi-scale system response of the sub-economy, that will further reduce the fluctuations in the value of the currency and stabilize the economy. The currency management firm will coordinate the fiscal investment policy with the monetary policy, issuing the currency needed for investment based on the projected demand that those investments will create. This will lead to a modest long-term stable growth in the value of the currency.
The reserves are essential to the stability of the economy. They give the dynamical financial system the inertia to change that stabilizes the economy. As shown in Fig. 4, the financial structure of the currency equity (fully backed security by tangible assets) and the stock equity (fully backed security by intangible assets) eliminates the super-charged risk of a highly leveraged derivative security as shown in Fig. 27. First of all, the liquid reserves (whether that be gold, cash or inventory like stored hydrocarbons) give the assets that can be sold by the currency management firm to purchase the electronic currency as directed by the genAI, to stabilize the value of the electronic currency. It is anticipated that less cash reserves are needed for a larger sub-economy since the relative size of the fluctuations are smaller. Then there is the longer-term strategic monetary policy associated with execution of the fiscal investment policy and the control of longer term inflation and deflation. In times of deflation (recession), this will involve issuance of new electronic currency to buy liquid assets, subsidize labor costs, buy capital assets, and invest in existing capital assets. This will position the currency management firm to respond to periods of inflation by selling the liquid assets, terminating the labor subsidies, selling capital assets and reducing the investment in existing capital assets. This buying and selling is in contrast to the loaning and borrowing of sovereign currencies. Most governments do not have either the liquid assets (even if they do have a gold reserve like the US had at one time they do not actively trade it, it is only there to be a physical standard), or the capital assets to sell in a time of inflation. They rely only on the issuance of debt. A notable exception is the recent sale of the US petroleum reserve to combat the post COVID inflation. There are additional ways that the reserves increase the resiliency that will be discussed in the following paragraphs of this section.
@@
Before we proceed to that phase, let us examine why the system that we are proposing does not have a “risk squeeze”. The reason is simple. This system does not make the local assumption and securitizes the loan using the real model of risk – a multi-scale model of risk using genAI.
@@
In summary, our system is protected from economic collapse by: (1) significant (about 10%) liquid reserves, (2) complete capital reserves (that is the electronic currency being used for transactions and working capital), (3) genAI and non-debt based methods of monetary control, with insurance, matched to the structure of the economy, (4) a floating currency value, and (5) the use of electronic currency leveraged, not debt leveraged financing.
XVII Existing currencies
Existing currencies are a grab bag of misfits that range from the mundane to fraud. Let’s start by examining the ubiquitous sovereign currency, that has been the only type of currency until recently. Historical currencies have been based on a precious metal (gold or silver in most cases) liquid reserve that has literally been attached as a standard to the coin as the material from which it was fabricated. In more recent history, paper notes were issued and pledged that there were gold or silver reserves in the vault of the sovereign. In the most recent history, sovereigns have floated the currency, eliminating the gold reserves and backing the currency with the tangible and intangible assets of the sovereignty, with negligible liquid reserves in many cases. Monetary policy has been controlled by the literal minting of coins and printing of currency, then having the central bank loan to commercial banks at a prime interest rate. To reduce the currency in circulation the sovereign issues debt (i.e., US T-bills and British GILTs). Rarely has the sovereign bought and sold liquid reserves as an method of implementing monetary policy. A recent exception is the selling of the US Petroleum Reserve by the US Government as a measure to combat the post-COVID supply-chain induced inflation. As discussed previously in Sec. XVI, this is a poor method of system control at best and is unstable in many cases. In addition, there is usually an uncoordinated fiscal investment policy that is driven by politics, not economics.